QUESTION 7

  1. A dipole consisting of charges ±e, 218 nm apart placed between two very large (essentially infinite) sheets carrying equal but opposite charge densities of 138 μm/m2. What is the maximum magnitude of the torque the sheets can exert of the dipole? (Give your answer in scientific notation using N.m as unit)
Answer #1

electric field between plates

E = sigma / ( eo)

E = 138*10^-6 / (8.85*10^-12)

E

now max torque is given by

T = eE (d/2)

T = 1.6*10^-19* 1.56*10^5* (218/2)*10^-9

T = 2.72*10^-21 Nm

=======

Comment in case any doubt, will reply for sure, goodluck